OpenStudy (samara8954):

Fred is playing tennis. For every 2 of his serves that land in, 5 serves land out. If he hit 26 serves in, how many serves landed out? 10 260 65 130

OpenStudy (samara8954):

@jacobwat

OpenStudy (samara8954):

@dtuniyants

OpenStudy (anonymous):

1 group=2 land in+5 land out. there are 26/2 groups. =13. So there are 13*5=65 of his serves land out.

OpenStudy (samara8954):

Ivan is an artist. He creates 1 sculpture for every 7 paintings. If he has created 42 paintings, how many sculptures has he created?

OpenStudy (anonymous):

26 divided by 2 then multiplied by 5

OpenStudy (samara8954):

35 6 294 42

OpenStudy (cwrw238):

for every 2 serves landing in 5 land out by proportion if 26 land in then (5 * 26) / 2 land out

OpenStudy (samara8954):

is the answer 6 to the new problem

OpenStudy (anonymous):

1 group =1 sc+7 pt. 42/7 =6 groups ---------->6*1=6 sculptures for 42 paintings

OpenStudy (samara8954):

Glenn and Helen are playing ping pong. For every 3 points that Glenn scores, Helen scores 4 points. If Glenn scores 9 points, how many points does Helen score? 36 12 7 108 END STUDY SESSION

OpenStudy (anonymous):

1 group =3 Glenn+4 Helen 9/3=3 groups---------->3*4=12 points will score Helen

OpenStudy (samara8954):

Daniel watches 1 hour of television for every 3 hours of homework he does each week. If he did 12 hours of homework last week, how many hours of television did he watch? 12 36 4 9

OpenStudy (samara8954):

is the answer 4

OpenStudy (anonymous):

yes

OpenStudy (samara8954):

Solve for n. n = 320 n = 44 n = 11 n = 4

OpenStudy (anonymous):

ok. this is about finding the equivalent fraction.

OpenStudy (anonymous):

if 4 goes into 32 8 times. then what goes into 88 8 times? @Samara8954

OpenStudy (anonymous):

you can solve this by multiplying each n value by 8 until you find one that equals 88